# Traceable Black-Box Watermarks For Federated Learning

### v3

@@ 3 Problem Formulation
4 $\tilde{\theta}_{i}\leftarrow\mathbf{M}_{m}\odot(1/n)\sum^{n}_{i=1}(\theta_{i}+\Delta_{i})+\mathbf{M}_{w}\odot(\theta_{i}+\Delta_{i})$
@@ 10 $\theta^{\prime}_{i}\leftarrow\tilde{\theta}^{\tau_{w}}_{i}$
12
13Return $\{\theta^{\prime}_{i}\}^{n}_{i=1}$
@@
4 Injecting Traceable Black-box Watermarks
Based on the insights in Remark 2, we propose a novel method called $\mathbf{TraMark}$ detailed in Algorithm 1, which can be easily integrated into existing FedAvg frameworks to solve Problem 1. We give the complete process of FedAvg with $\mathbf{TraMark}$ in Algorithm 2 in the Appendix A.3. Specifically, $\mathbf{TraMark}$ operates entirely on the server side. In each training round, once the server receives local model updates from the clients, it uses $\mathbf{TraMark}$ to aggregate these updates and derive a personalized global model for each client. For each global model, the watermark is injected by learning from a distinct watermark dataset. The server then sends the watermarked global model back to each client for the next round of training or deployment.
4.1 Constraining Watermarking Region
Existing watermarking approaches either retrain the global model directly on the watermark dataset (Tekgul et al. 2021; Shao et al. 2024) or require local clients to collaboratively inject watermarks (Li et al. 2022; Liu et al. 2021; Nie and Lu 2024b). However, these methods cause watermark-related perturbations to spread across the entire parameter space, leading to two key issues. First, the dispersed watermark perturbations may significantly degrade the main task performance. Second, even if each client’s model embeds a unique watermark, model aggregation in the next round fuses these watermarks, causing collisions that compromise traceability. To mitigate the impact on main task performance and ensure traceability, $\mathbf{TraMark}$ restricts watermarking to a small subset of the model’s parameter space. Only this designated watermarking region carries the watermark, and its parameters are excluded from model aggregation, preserving distinct watermarks for each client in the next training round. Specifically, in $\mathbf{TraMark}$, given a model $\theta\in\mathbb{R}^{d}$, the server partitions the whole parameter space into watermarking region and main task region with a partition ratio $k\in[0,1)$, resulting in two complementary binary masks:
  • The watermarking mask $\mathbf{M}_{w}\in\{0,1\}^{d}$, where $\mathbf{M}_{w}]_{j}=1$ means that the $j$-th parameter in $\theta$ is used for watermarking task and $\mathrm{sum}(\mathbf{M}_{w})=k\times d$.
  • The main task mask $\mathbf{M}_{m}\in\{0,1\}^{d}$, where $\mathbf{M}_{m}]_{k}=1$ means that the $k$-th parameter in $\theta$ is used for main task and $\mathrm{sum}(\mathbf{M}_{m})=(1-k)\times d$.
These two complementary masks ensure that all model parameters are fully partitioned into the watermarking and main task regions (i.e., $\mathbf{M}_{w}+\mathbf{M}_{m}=\mathbf{1}^{d}$). Moreover, once determined, the masks remain unchanged throughout the entire watermarking process. We will discuss how to partition the model later in Section 4.4.
@@ 4.2 Masked Aggregation & Watermark Injection
Masked Aggregation.
With the constrained watermarking region, $\mathbf{TraMark}$ leverages a novel masked aggregation method to avoid watermark collision. Specifically, instead of applying a naive aggregation approach (e.g., FedAvg), $\mathbf{TraMark}$ aggregates the parameters in the main task region only and prevents the watermarking region from parameter fusion. In detail, in each training round, given the model updates $\{\Delta_{i}\}_{i=1}^{n}$, the server aggregates them to generate the personalized global model for each client individually via $\tilde{\theta}_{i}=\mathbf{M}_{m}\odot(1/n)\sum^{n}_{i=1}(\theta_{i}+\Delta_{i})+\mathbf{M}_{w}\odot(\theta_{i}+\Delta_{i}),\ \forall i\in[n]$ (Line 1 in Algorithm 1). Here, the first term $\mathbf{M}_{m}\odot(1/n)\sum^{n}_{i=1}(\theta_{i}+\Delta_{i})$ averages the model updates in the main task region, and the second term $\mathbf{M}_{w}\odot(\theta_{i}+\Delta_{i})$ preserves the update of client $i$ in the watermarking region. In this case, the server generates a personalized global model for each client, which always contains a distinct watermark for the client while still benefiting from the aggregated model updates for the main task.
Watermark Injection.
For each personalized global model $\tilde{\theta}_{i},\forall i\in[n]$, its watermark is injected by training $\tilde{\theta}_{i}$ on the corresponding distinct watermark dataset $\mathcal{D}^{w}_{i}$ for $\tau_{w}$ iterations. In this process, only the watermarking region is updated, ensuring that knowledge from the watermark dataset does not spread to the main task region (Line 11). Technically, in each step of local training, the mini-batch gradient is multiplied by $\mathbf{M}_{w}$ (Line 11), zeroing out the gradients for the main task region to avoid the impact of the watermark on the main task. After watermark injection, the server obtains the watermarked global models, which will be sent to the clients to perform their main tasks of the next training round or deployment (Line 1). Since the clients’ local training protocol for the main task remains unchanged, model updates continue across the entire parameter space. Notably, this case leads to a potential risk: over time, the embedded watermarks may gradually fade. To prevent watermark fading, $\mathbf{TraMark}$ can be applied at every training round to continuously enhance the watermark, as advised by prior works (Tekgul et al. 2021; Shao et al. 2024; Nie and Lu 2024b; Li et al. 2022).
@@ 4.3 Distinct Watermark Dataset
As noted in Remark 1, ensuring sufficient differences between watermark datasets is crucial for learning distinct watermarks and preventing collisions. To achieve this, in the $\mathbf{TraMark}$ , the server assigns each global model a unique watermark dataset. Specifically, the watermark datasets designed for different models should differ from each other in both their triggers and their output distribution. Let $\mathcal{D}^{w}_{i}=\{(x,\phi_{i}(x))\mid x\in\mathcal{X}^{w}_{i}\}$ denote the watermark dataset for personalized global model $\tilde{\theta}_{i}$, where $\mathcal{X}^{w}_{i}\cap\mathcal{X}^{w}_{j}=\emptyset$ for any $i\neq j,\ \forall i,j\in[n]$. Furthermore, each client is assigned a unique output distribution $\phi_{i}(x)$, guaranteeing that $\phi_{i}(x)\neq\phi_{j}(x)$. For trigger selection, existing methods have explored various approaches, including randomly generated patterns (Tekgul et al. 2021; Shao et al. 2024), adversarially perturbed samples (Li et al. 2022), and samples embedded with backdoor triggers (Liu et al. 2021). In our case, to ensure each client receives a maximally distinct trigger, we select out-of-distribution samples absent from the main task dataset. This guarantees that the learned watermark remains independent of the main task. For example, in a classifier trained for traffic sign recognition, per-label samples from the MNIST dataset serve as effective triggers for different clients. Assigning a distinct watermark dataset to each personalized global model ensures that a watermarked model responds only to triggers from its own dataset, mapping them to the predefined output. When exposed to triggers from other watermark datasets, it produces random guesses, effectively minimizing the risk of collisions. This distinct watermark dataset assignment may potentially limit the scalability of $\mathbf{TraMark}$ , particularly when the learning task involves only a few labels. We discuss how to address this in Appendix A.10.
@@ 4.4 Selection of Watermarking Region
A key question that remains is: how should the server select the watermarking region? A naive approach is to randomly assign a small portion of parameters for watermark injection before training begins. However, this risks degrading main task performance, as critical parameters for the main task may be allocated to the watermarking region, leading to main task performance loss. Recall from Remark 2 that maximizing $\|\delta_{i}-\delta_{j}\|^{2}_{2}$ is crucial for avoiding collisions. However, excessive divergence between $\delta_{i}$ and $\delta_{j}$ may negatively impact main task performance. Given that $\|\delta_{i}-\delta_{j}\|^{2}_{2}=\|\mathbf{M}_{m}\odot(\delta_{i}-\delta_{j})\|^{2}_{2}+\|\mathbf{M}_{w}\odot(\delta_{i}-\delta_{j})\|^{2}_{2},$ where $\mathbf{M}_{m}=\mathbf{1}^{d}-\mathbf{M}_{w}$ and the watermarking process is confined to the watermarking region, the objective simplifies to maximizing $\|\mathbf{M}_{w}\odot(\delta_{i}-\delta_{j})\|^{2}_{2}$. This ensures that watermark injection and collision avoidance should not affect parameters in the main task region. Consequently, if $\mathbf{M}_{m}$ contains the most important parameters while $\mathbf{M}_{w}$ is assigned to unimportant ones, the primary accuracy remains largely unaffected. Typically, parameter importance is measured by magnitude (absolute value), with larger values indicating greater importance (Xu, Zhang, and Hu 2025; Hu, Guo, and Gong 2023; Panda et al. 2022). However, since network parameters are randomly initialized at the start of training, their importance is not yet established. As a result, assigning parameters to regions too early may lead to suboptimal partitioning, potentially degrading main task performance. To this end, $\mathbf{TraMark}$ introduces a warmup training phase, where the global model undergoes standard federated training (e.g., FedAvg) for $\alpha\times T$ rounds before watermarking. The warmup training ratio $\alpha\in[0,1)$ determines the fraction of total training rounds allocated to this phase, ensuring the model is robust enough to the main task before starting the watermark injection. Once warmup training is complete, the server obtains the watermarking region by selecting $k\times d$} least important parameters (i.e., those that have the smallest absolute values), and the remaining parameters are assigned to the main task region. We present the detailed technical procedure for selecting the watermarking region in Algorithm 2, included in Appendix A.3.
@@ 5.1 Experimental Settings
Following previous works (Shao et al. 2024; Zhang et al. 2024; Nie and Lu 2024a, b), we evaluate our methods on FMNIST (Xiao, Rasul, and Vollgraf 2017), CIFAR-10 (Krizhevsky et al. 2009), and CIFAR-100 (Krizhevsky et al. 2009), using a CNN, AlexNet (Krizhevsky, Sutskever, and Hinton 2012), and VGG-16 (Simonyan 2014), respectively. In addition, we test $\mathbf{TraMark}$ on large-scale Tiny-ImageNet using ViT (Dosovitskiy 2020). For all datasets, we consider both independent and identically distributed (IID) data and non-IID data scenarios. To simulate non-IID cases, we use the Dirichlet distribution (Minka 2000) with a default degree $\gamma=0.5$. Following previous works (Li et al. 2022; Xu et al. 2024; Nie and Lu 2024a, b), we set up a cross-silo FL system with $10$ clients. We also test $\mathbf{TraMark}$ on large-scale FL with client sampling in Appendix A.8. Each client performs local training with $\tau_{l}=5$ iterations and a learning rate of $\eta_{l}=0.01$. The training rounds for FMNIST, CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets are $50,100,100,$ and $50$. All experiments are repeated $3$ times with different seeds and we report the averaged results.
Baselines and $\mathbf{TraMark}$ Settings.
@@
In all experiments, we use the MNIST (LeCun et al. 1998) dataset as the source for watermarking. Each global model is assigned a watermark dataset containing samples from a distinct MNIST label, ensuring both $\mathcal{X}^{w}_{i}\cap\mathcal{X}^{w}_{j}=\emptyset$ and $\phi_{i}(x)\neq\phi_{j}(x)$ for any two clients $i$ and $j$. We present the results of $\mathbf{TraMark}$ using other watermark datasets in Appendix A.7. Each watermark dataset consists of $100$ samples. The watermarking learning rate is set to $\eta_{w}=1e^{-4}$, and the number of watermarking iterations is $\tau_{w}=5$. The partition ratio $k$ is set to $1\%$. The warmup training ratio $\alpha$ is set to $0.5$. To ensure a fair comparison, we evaluate $\mathbf{TraMark}$ against two server-side watermarking approaches: WAFFLE (Tekgul et al. 2021), a black-box watermarking method that does not ensure traceability, and FedTracker (Shao et al. 2024), a white-box watermarking method that ensures traceability.
Evaluation Metrics.
@@
We evaluate the performance of each method using two key metrics: main task accuracy (MA) and model leakage verification rate (VR). MA is measured using the main task test set. Since $\mathbf{TraMark}$ and FedTracker introduce slight variations in each local model due to watermark injection, we compute MA as the average accuracy across all local models, following (Shao et al. 2024). VR quantifies the proportion of watermarked models that are successfully attributed to their respective owners. We evaluate each watermarked model on the full test set, compute the per-label accuracy, and identify the label with the highest accuracy. If this highest-accuracy label matches the pre-assigned label of the model’s owner, the model is considered successfully verified (more details are given in Appendix A.4). For FedTracker, we follow its original definition of VR, where traceability is determined based on fingerprint similarity in a white-box setting.
5.2 Empirical Results
Figure 1: Verification confidence and verification leakage across training rounds on CIFAR-10 and CIFAR-100.
Verification Interval.
@@ 5.2 Empirical Results
We first demonstrate the verifier’s confidence in identifying the leaker of suspect models embedded with watermarks injected by $\mathbf{TraMark}$. Specifically, we calculate two key metrics: verification confidence, the test accuracy of a watermarked model on its own watermarking dataset, and verification leakage, its average test accuracy on other clients’ watermarking datasets. The difference between these two metrics, termed the verification interval, reflects the verifier’s confidence. Figure 1 illustrates the averaged verification confidence and leakage across training rounds on CIFAR-10 and CIFAR-100 datasets. We observe that $\mathbf{TraMark}$ maintains a consistently large verification interval throughout training. Moreover, as training progresses, the interval widens due to a steady increase in verification confidence, indicating that watermark injection in $\mathbf{TraMark}$ continuously enhances watermark effectiveness despite potential performance degradation from local training. Additionally, since $\mathbf{TraMark}$ injects watermarks only within the designated watermarking region and employs masked aggregation, the watermarked model consistently performs poorly on other clients’ watermarking datasets. These factors collectively contribute to a clear verification interval, ensuring successful model leakage verification. Further results on the changes in the divergence of each watermarked model’s output on its watermark dataset (the constraint in Problem 1) are provided in Appendix A.6.
@@
Table 1: Comprehensive comparison of MA and VR between different methods under both IID and non-IID settings (highlighted with a “N”). MAs are shown in percentages ($\%$). If a method achieves a satisfactory VR (exceeds $95\%$), we denote it with “✓”; otherwise, we use “✗”.
@@
FM $92.60$ - $92.21$ $89.95$ $91.20$
FM (N) $91.52$ - $91.41$ $67.50$ $91.31$
C-10 $89.15$ - $89.16$ $87.56$ $88.58$
C-10 (N) $87.01$ - $86.75$ $83.42$ $86.26$
C-100 $61.91$ - $61.68$ $61.05$ $61.13$
C-100 (N) $60.19$ - $60.04$ $60.12$ $58.95$
Tiny $21.05$ - $21.24$ $20.40$ $20.91$
Tiny (N) $20.09$ - $19.97$ $20.00$ $20.06$
Average $65.44$ - $65.31$ $61.25$ $64.90$
Main Results.
@@
We report the MA and VR of each method on FMNIST (FM), CIFAR-10 (C-10), CIFAR-100 (C-100), and Tiny-ImageNet (Tiny) under both IID and non-IID settings in Table 1. Overall, $\mathbf{TraMark}$ effectively injects traceable watermarks into personalized global models while preserving high model performance. It consistently achieves a high VR across all datasets, maintaining an average of $99.17\%$. In contrast, FedTracker fails to ensure satisfactory traceability on CIFAR-10, resulting in an average VR of only $87.50\%$. This instability is due to the injection of key matrices into model parameters and the absence of explicit mechanisms to prevent watermark collisions after aggregation. Regarding MA, $\mathbf{TraMark}$ exhibits strong model performance, with only a $0.54\%$ drop compared to FedAvg. While WAFFLE achieves a slightly higher average MA ($0.41\%$ above $\mathbf{TraMark}$), it does not guarantee the traceability of watermarked models. FedTracker, on the other hand, suffers a significant $4.19\%$ decline in MA due to the unconstrained watermarking region, which compromises model utility. In conclusion, $\mathbf{TraMark}$ successfully embeds traceable black-box watermarks while incurring minimal performance loss, making it a robust and practical watermarking solution for FL.
@@
Figure 2: Averaged MA and VR results of $\mathbf{TraMark}$ under pruning and fine-tuning attacks across four datasets.
Robustness to Attacks.
@@
We evaluate the robustness of watermarked models trained by $\mathbf{TraMark}$ against pruning and fine-tuning attacks. Specifically, malicious clients may prune or fine-tune their local models to remove or reduce the effectiveness of watermarks. For the pruning attack, we test pruning ratios ranging from $30\%$ to an extreme $99\%$. For fine-tuning attacks, we assume that malicious clients fine-tune their models on their local datasets for $30$ epochs, using the same learning rate as in the original training process. The averaged MA and VR results of $\mathbf{TraMark}$ across four datasets are summarized in Figure 2. With moderate pruning ratios ($30\%$ to $70\%$), MA remains largely unaffected, while VR is also preserved. As the pruning ratio increases, MA declines rapidly, accompanied by a decrease in VR. These results demonstrate that the parameters in the watermarking region are coupled with the main task parameters, making simple magnitude-based pruning ineffective in removing the watermarks. This coupling also contributes to stable VR across various fine-tuning epochs during the fine-tuning attack. Additionally, we evaluate $\mathbf{TraMark}$ against fine-tuning attack with various fine-tuning learning rates, the quantization attack, and stronger adaptive attack, with results provided in Appendix A.9.

### v4

@@ 4.1 Constraining Watermarking Region
3
4 $\tilde{\theta}_{i}\leftarrow\theta_{i}+\mathbf{M}_{m}\odot\frac{1}{n}\sum^{n}_{i=1}\Delta_{i}+\mathbf{M}_{w}\odot\Delta_{i}$
@@ 11 $\theta^{\prime}_{i}\leftarrow\tilde{\theta}^{\tau_{w}}_{i}$
Return $\{\theta^{\prime}_{i}\}^{n}_{i=1}$
@@
Existing watermarking approaches either retrain the global model directly on the watermark dataset (Tekgul et al., 2021; Shao et al., 2024) or require local clients to collaboratively inject watermarks (Li et al., 2022a; Liu et al., 2021; Nie and Lu, 2024b). However, these methods cause watermark-related perturbations to spread across the entire parameter space, leading to two key issues. First, the dispersed watermark perturbations may significantly degrade the main task performance. Second, even if each client’s model embeds a unique watermark, model aggregation in the next round fuses these watermarks, causing collisions that compromise traceability. To mitigate the impact on main task performance and ensure traceability, $\mathbf{TraMark}$ restricts watermarking to a small subset of the model’s parameter space. Only this designated watermarking region carries the watermark, and its parameters are excluded from model aggregation, preserving distinct watermarks for each client in the next training round. Specifically, in $\mathbf{TraMark}$, given a model $\theta\in\mathbb{R}^{d}$, it partitions the whole parameter space into watermarking region and main task region with a partition ratio $k\in[0,1)$, resulting in two complementary binary masks:
  • The watermarking mask $\mathbf{M}_{w}\in\{0,1\}^{d}$, where $\mathbf{M}_{w}]_{j}=1$ means that the $j$-th parameter in $\theta$ is used for watermarking task and $\mathtt{sum}(\mathbf{M}_{w})=k\times d$.
  • The main task mask $\mathbf{M}_{m}\in\{0,1\}^{d}$, where $\mathbf{M}_{m}]_{p}=1$ means that the $p$-th parameter in $\theta$ is used for main task and $\mathtt{sum}(\mathbf{M}_{m})=(1-k)\times d$.
These two complementary masks (i.e., $\mathbf{M}_{w}+\mathbf{M}_{m}=\mathbf{1}^{d}$) ensure that all model parameters in the model $\theta$ are fully partitioned into the watermarking and main task regions. Moreover, once determined, the masks remain unchanged throughout the entire watermarking process. We will discuss how to get these two masks later in Section 4.4.
@@ 4.2 Masked Aggregation & Watermark Injection
Masked Aggregation. With the constrained watermarking region, $\mathbf{TraMark}$ leverages a novel masked aggregation method to avoid watermark collision. Specifically, instead of applying a naive aggregation approach (e.g., FedAvg), $\mathbf{TraMark}$ aggregates parameters only in the main task region and prevents the watermarking region from parameter fusion. In detail, in each training round, given the model updates $\{\Delta_{i}\}_{i=1}^{n}$, the server aggregates them to generate the personalized global model for each client individually via $\tilde{\theta}_{i}=\theta_{i}+\mathbf{M}_{m}\odot\frac{1}{n}\sum^{n}_{i=1}\Delta_{i}+\mathbf{M}_{w}\odot\Delta_{i},\ \forall i\in[n]$ (Line 1 in Algorithm 1). Here, the second term $\mathbf{M}_{m}\odot\frac{1}{n}\sum^{n}_{i=1}\Delta_{i}$ averages the model updates in the main task region, and the third term $\mathbf{M}_{w}\odot\Delta_{i}$ preserves the parameters in the watermarking region for the model update of client $i$. In this case, the server generates a personalized global model for each client, which always contains a distinct watermark for the client while still benefiting from the aggregated model updates for the main task.
Watermark Injection. For each personalized global model $\tilde{\theta}_{i},\forall i\in[n]$, its watermark is injected by training $\tilde{\theta}_{i}$ on the corresponding distinct watermark dataset $\mathcal{D}^{w}_{i}$ for $\tau_{w}$ iterations. In this process, only the watermarking region is updated, ensuring that knowledge from the watermark dataset does not spread to the main task region (Line 11). Technically, in each step of local training, the mini-batch gradient $g^{s}_{i}$ is multiplied by $\mathbf{M}_{w}$ (Line 11), zeroing out the gradients for the main task region to avoid the impact of the watermark on the main task. After watermark injection, the server obtains the watermarked global models, which will be sent to the clients to perform their main tasks of the next training round or deployment (Line 1). Since the clients’ local training protocol for the main task remains unchanged, model updates continue across the entire parameter space. However, this introduces a potential risk: the embedded watermarks may gradually fade over time. To this end, $\mathbf{TraMark}$ can be applied at every training round to continuously preserve the watermark, as advised by prior works (Tekgul et al., 2021; Shao et al., 2024; Nie and Lu, 2024b; Li et al., 2022a).
@@ 4.3 Distinct Watermark Dataset
As noted in Remark 1, ensuring sufficient dissimilarity between watermark datasets is crucial for learning distinct watermarks and preventing collisions. To achieve this, in the $\mathbf{TraMark}$, the server assigns each global model a unique watermark dataset. Specifically, the watermark datasets designed for different models should differ from each other in both their triggers and their output distributions. Let $\mathcal{D}^{w}_{i}=\{(x,\phi_{i}(x))\mid x\in\mathcal{X}^{w}_{i}\}$ denote the watermark dataset for personalized global model $\tilde{\theta}_{i}$, where $\mathcal{X}^{w}_{i}\cap\mathcal{X}^{w}_{j}=\emptyset$ for any $i\neq j,\ \forall i,j\in[n]$. Furthermore, each client is assigned a unique output distribution $\phi_{i}(x)$, guaranteeing that $\phi_{i}(x)\neq\phi_{j}(x)$. For trigger selection, existing methods have explored various approaches, including randomly generated patterns (Tekgul et al., 2021; Shao et al., 2024), adversarially perturbed samples (Li et al., 2022a), and samples embedded with backdoor triggers (Liu et al., 2021). In our case, to ensure each client receives a maximally distinct trigger, we select out-of-distribution samples absent from the main task dataset. This guarantees that the learned watermark remains independent of the main task. For example, in a classifier trained for traffic sign recognition, per-label samples from the MNIST dataset serve as effective triggers for different clients. Assigning a distinct watermark dataset to each personalized global model ensures that a watermarked model responds only to triggers from its own dataset, mapping them to the predefined output. When exposed to triggers from other watermark datasets, it produces random guesses, effectively minimizing the risk of collisions. This distinct watermark dataset assignment may limit the scalability of $\mathbf{TraMark}$, particularly in learning tasks with only a few labels. We discuss possible solutions to this limitation in Appendix A.9.
@@ 4.4 Selection of Watermarking Region
Recall from Remark 2 that maximizing $\|\delta_{i}-\delta_{j}\|^{2}_{2}$ is crucial for avoiding collisions. However, excessive divergence between $\delta_{i}$ and $\delta_{j}$ may negatively impact main task performance. Given that $\|\delta_{i}-\delta_{j}\|^{2}_{2}=\|\mathbf{M}_{m}\odot(\delta_{i}-\delta_{j})\|^{2}_{2}+\|\mathbf{M}_{w}\odot(\delta_{i}-\delta_{j})\|^{2}_{2},$ where $\mathbf{M}_{m}=\mathbf{1}^{d}-\mathbf{M}_{w}$ and the watermarking process is confined to the watermarking region, the objective simplifies to maximizing $\|\mathbf{M}_{w}\odot(\delta_{i}-\delta_{j})\|^{2}_{2}$. This ensures that watermark injection and collision avoidance should not affect parameters in the main task region. Consequently, if $\mathbf{M}_{m}$ contains the most important parameters while $\mathbf{M}_{w}$ is assigned to unimportant ones, the primary accuracy remains largely unaffected. Typically, parameter importance is measured by magnitude (absolute value), with larger values indicating greater importance (Xu et al., 2025; Hu et al., 2023; Panda et al., 2022). However, since network parameters are randomly initialized at the start of training, their importance is not yet established. As a result, assigning parameters to regions too early may lead to suboptimal partitioning, potentially degrading main task performance. To this end, $\mathbf{TraMark}$ introduces a warmup training phase, where the global model undergoes standard federated training (e.g., FedAvg) for $\alpha\times T$ rounds before watermarking. The warmup training ratio $\alpha\in[0,1)$ determines the fraction of total training rounds allocated to this phase, ensuring the model is robust enough to the main task before starting the watermark injection. Once warmup training is complete, the server obtains the watermarking region by selecting $k\times d$} least important parameters (i.e., those that have the smallest absolute values), and the remaining parameters are assigned to the main task region. We present the detailed technical procedure of selecting the watermarking region in Algorithm 2, included in Appendix A.3.
@@ 5.1 Experimental Settings
General Settings. Following (Shao et al., 2024; Zhang et al., 2024; Nie and Lu, 2024a; b), we evaluate our methods on FMNIST (Xiao et al., 2017), CIFAR-10 (Krizhevsky and others, 2009), and CIFAR-100 (Krizhevsky and others, 2009), using a CNN, AlexNet (Krizhevsky et al., 2012), and VGG-16 (Simonyan, 2014), respectively. Besides these benchmarks, we test $\mathbf{TraMark}$ on large-scale Tiny-ImageNet using ViT (Dosovitskiy et al., 2021). For all datasets, we consider both independent and identically distributed (IID) data and non-IID data scenarios. To simulate non-IID cases, we use the Dirichlet distribution (Minka, 2000) with a default degree $\gamma=0.5$. Following previous works (Li et al., 2022a; Xu et al., 2024; Nie and Lu, 2024a; b), we set up a cross-silo FL system with $10$ clients. We also test $\mathbf{TraMark}$ on larger-scale $n=50$ FL with client sampling in Appendix A.7. Each client performs local training with $\tau_{l}=5$ iterations and a learning rate of $\eta_{l}=0.01$. The training rounds for FMNIST, CIFAR-10, CIFAR-100, and Tiny-ImageNet are $50,100,100,$ and $50$. All experiments are repeated $3$ times with different seeds, and we report the averaged results.
@@
Baselines and $\mathbf{TraMark}$ Settings. In all experiments, we use the MNIST (LeCun et al., 1998) dataset as the source for watermarking. Each global model is assigned a watermark dataset containing samples from a distinct MNIST label, ensuring both $\mathcal{X}^{w}_{i}\cap\mathcal{X}^{w}_{j}=\emptyset$ and $\phi_{i}(x)\neq\phi_{j}(x)$ for any two clients $i$ and $j$. Furthermore, we demonstrate that $\mathbf{TraMark}$ can accommodate various types of watermark datasets, including randomly generated patterns as proposed in (Tekgul et al., 2021), with results provided in Section 5.3. Each watermark dataset consists of $100$ samples. The watermarking learning rate is set to $\eta_{w}=1e^{-4}$, and the number of watermarking iterations is $\tau_{w}=5$. The partition ratio $k$ is set to $1\%$. The warmup training ratio $\alpha$ is set to $0.5$. To ensure a fair comparison, we evaluate $\mathbf{TraMark}$ against two server-side watermarking approaches: WAFFLE (Tekgul et al., 2021), a black-box watermarking method that does not ensure traceability, and FedTracker (Shao et al., 2024), a white-box watermarking method that ensures traceability.
@@
Evaluation Metrics. We evaluate the performance of each method using two key metrics: main task accuracy (MA) and model leakage verification rate (VR). MA is measured using the main task test set. Since $\mathbf{TraMark}$ and FedTracker introduce slight variations in each local model due to watermark injection, we compute MA as the average accuracy across all local models, following (Shao et al., 2024). VR quantifies the proportion of watermarked models that are successfully attributed to their respective owners. We evaluate each watermarked model on the full test set, compute the per-label accuracy, and identify the label with the highest accuracy. If this highest-accuracy label matches the pre-assigned label of the model’s owner, the model is considered successfully verified (more details are given in Appendix A.4). For FedTracker, we follow its original definition of VR, where traceability is determined based on fingerprint similarity in a white-box setting.
5.2 Main Empirical Results
@@
Verification Interval. We first demonstrate the verifier’s confidence in identifying the leaker of suspect models embedded with watermarks injected by $\mathbf{TraMark}$. Specifically, we compute two key metrics: (1) the test accuracy of each watermarked model on its own watermarking dataset, referred to as verification confidence; and (2) the average test accuracy of the model on other clients’ watermarking datasets, referred to as verification leakage. The difference between these two metrics, termed the verification interval, reflects the verifier’s confidence. Figure 1 illustrates the averaged verification confidence, leakage, interval, and the VR changes across training rounds on CIFAR-10 and CIFAR-100 datasets.
Figure 1: Verification (Ver) confidence, verification leakage, verification interval, and VR on CIFAR-10 and CIFAR-100.
We observe that as training progresses, the verification interval widens due to a steady increase in verification confidence. Additionally, since $\mathbf{TraMark}$ injects watermarks only within the designated watermarking region and employs masked aggregation, the watermarked model consistently performs poorly on other clients’ watermarking datasets. These factors together form a clear verification interval, which ensures reliable model leakage verification. This is supported by the high VR obtained by $\mathbf{TraMark}$. Further results on the changes in the divergence of each watermarked model’s output on its watermark dataset (the constraint in Problem 1) are provided in Appendix A.6.
@@
Table 1: Comprehensive comparison of MA and VR between different methods under both IID and non-IID settings (highlighted with a “N”). MAs and VRs are shown in percentages ($\%$).
@@
FM $92.60$ - $92.21$ - $89.95$ $100.00$ $91.20$ $96.67$
FM (N) $91.52$ - $91.41$ - $67.50$ $100.00$ $91.31$ $100.00$
C-10 $89.15$ - $89.16$ - $87.56$ $60.00$ $88.58$ $100.00$
C-10 (N) $87.01$ - $86.75$ - $83.42$ $50.00$ $86.26$ $100.00$
C-100 $61.91$ - $61.68$ - $61.05$ $100.00$ $61.13$ $100.00$
C-100 (N) $60.19$ - $60.04$ - $60.12$ $90.00$ $58.95$ $100.00$
Tiny $21.05$ - $21.24$ - $20.40$ $100.00$ $20.91$ $100.00$
Tiny (N) $20.09$ - $19.97$ - $20.00$ $100.00$ $20.06$ $100.00$
Average $65.44$ - $65.31$ - $61.25$ $87.50$ $64.90$ $99.58$
@@
Main Results. We report the MA and VR of each method on FMNIST (FM), CIFAR-10 (C-10), CIFAR-100 (C-100), and Tiny-ImageNet (Tiny) under both IID and non-IID settings in Table 1. Overall, $\mathbf{TraMark}$ effectively injects traceable watermarks into personalized global models while preserving high model performance. It consistently achieves a high VR across all datasets, maintaining an average of $99.58\%$. In contrast, FedTracker fails to ensure satisfactory traceability on CIFAR-10, resulting in an average VR of only $87.50\%$. This instability is due to the injection of key matrices into model parameters and the absence of explicit mechanisms to prevent watermark collisions after aggregation. Regarding MA, $\mathbf{TraMark}$ exhibits strong model performance, with only a $0.54\%$ drop compared to FedAvg. While WAFFLE achieves a slightly higher average MA ($0.41\%$ above $\mathbf{TraMark}$), it does not guarantee the traceability of watermarked models. FedTracker, on the other hand, suffers a significant $4.19\%$ decline in MA due to the unconstrained watermarking region, which compromises model utility. In conclusion, $\mathbf{TraMark}$ successfully embeds traceable black-box watermarks while incurring minimal performance loss, making it a robust and practical watermarking solution for FL.
@@
Figure 2: Averaged MA and VR results of $\mathbf{TraMark}$ under pruning and fine-tuning attacks.
@@
Robustness to Attacks. We evaluate the robustness of watermarked models trained by $\mathbf{TraMark}$ against pruning and fine-tuning attacks. Specifically, malicious clients may prune or fine-tune their local models to remove or reduce the effectiveness of watermarks. For the pruning attack, we test pruning ratios ranging from $30\%$ to an extreme $99\%$. For fine-tuning attacks, we assume that malicious clients fine-tune their models on their local datasets for $30$ epochs, using the same learning rate as in the original training process. The averaged MA and VR results of $\mathbf{TraMark}$ across four datasets are summarized in Figure 2. With moderate pruning ratios ($30\%$ to $70\%$), MA remains largely unaffected, while VR is also preserved. As the pruning ratio increases, MA declines rapidly, accompanied by a decrease in VR. These results demonstrate that the parameters in the watermarking region are coupled with the main task parameters, making simple magnitude-based pruning ineffective in removing the watermarks. This coupling also contributes to stable VR across various fine-tuning epochs. Additionally, we evaluate $\mathbf{TraMark}$ against fine-tuning attacks with different learning rates, quantization attacks, and the stronger adaptive RNP attack (Li et al., 2023) in Appendix A.8. The extended results further show the robustness of $\mathbf{TraMark}$.
5.3 Detailed Studies of $\mathbf{TraMark}$